News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 12 January 2001 12.00pm EST

During the last week, Chandra completed the observing schedule as planned.

Real-time procedures were conducted on Jan 10 to insert the LETG and on Jan 12 to retract the LETG, and a routine Ephemeris update was completed on Jan 9. The observations of Venus with ACIS-S and the LETG were completed successfully on Jan 9, however only 2 guide stars were tracked by the aspect camera during the observations due to higher than predicted light levels from the planet flooding the camera CCD chip. In order to reduce the risk of a transition to bright star hold during the 3rd Venus observation on Jan 13, a new star catalog was developed and the last segments of the week's load rebuilt and uplinked to the spacecraft. The observation will be monitoring in real-time.

Items of note from last week include 3 press releases from the AAS highlighting Chandra Science. The discovery of a glowing bubble of hot gas and an unexpected X-ray bright central star within the planetary nebula known as the Cat's Eye, the discovery that a known pulsar is the present-day counterpart to the supernova G11.2-0.3 that exploded in 386 AD, and the joint discovery with Hubble of the best direct evidence yet for the existence of an event horizon around a black hole. See for further details.

The planned observations for the coming week are shown below and include 400 ks observing the target H1821+643 following the insertion of the LETG on Jan 17.

Radiation Belts                      Jan 15 
1RXSJ0245+2620          HRC-S
HR 750                  HRC-I
2MASSJ000703+1554       ACIS-I
2MASSJ234449+1221       ACIS-I       Jan 16
3C 47                   ACIS-S
IC 1727                 ACIS-S
A262                    ACIS-S
NGC266                  ACIS-S       Jan 17
H1821+643               ACIS-S/LETG                     
Radiation Belts                      Jan 18
H1821+643               ACIS-S/LETG 
Radiation Belts                      Jan 20
H1821+643               ACIS-S/LETG  Jan 21 
Radiation Belts                      Jan 23


All spacecraft subsystems continued to operate nominally.

Return to Status Reports